How many license plates can be made using either two or three letters followed by either two or three digits?


1
Expert's answer
2022-01-05T04:04:02-0500

Number of license plates having either two or three letters followed by either two or three numbers:

Case 1: 2 letters followed by 2 numbers are "26^{2} \times 10^{2}"

Case 2: 2 letters followed by 3 numbers are "26^{2} \times 10^{3}"

Case 3: 3 letters followed by 2 numbers are "26^{3} \times 10^{2}"

Case 4: 3 letters followed by 3 numbers are "26^{3} \times 10^{3}"

So, the total number of license plates having two or three letters followed by two or three numbers

"=\left(26^{2} \times 10^{2}\right)+\left(26^{2} \times 10^{3}\right)+\left(26^{3} \times 10^{2}\right)+\left(26^{3} \times 10^{3}\right)=20,077,200"